North Eastern Locomotive Preservation Group

  • Increase font size
  • Default font size
  • Decrease font size
K1 Blog

A blog of activities associated with the K1 62005.



Friday 9th April 2021

Print

The K1 completed its work on the NYMR on Wednesday afternoon. The loco generally performed well but there are a few items to attend to before it leaves for Fort William (currently scheduled for 19th April). The most significant job is to replace the RH leading side rod bush. The bush currently fitted is new but it is too short so must be replaced. We will be taking the RH rods off and pushing the existing bush out tomorrow and the NYMR should complete the new bush by Wednesday at the latest.

Work on the new coach lighting, battery box modifications and kitchen redecoration has progressed but there is still much to do.

Many thanks to Mike B, Scott M, Richie Worrall, Angie, Rich P, Rich De, Sammy T, Rolly, Kenny M, Chris H, JO and Les Harper who have all worked on the engine or coach recently.

 

Sunday 4th April 2021

Print

Rolly, Angie, Kenny M, Mike B, new recruit Sammy Tidy and I worked on the loco and coach today. We changed a broken loco spring, prepared the engine for its next scheduled work (see below) and lit it up. On the coach a start was made on the redecoration of the kitchen and the alterations to one of the battery boxes was very nearly completed.

The K1 will be working 4 Grosmont/Whitby return trips a day on Monday, Tuesday and Wednesday. There will be working parties pretty much every day until the loco and coach depart for Fort William on 19th April.

 

Thursday 1st April 2021

Print

The K1 has been operating between Grosmont and Battersby since Monday. The only problem with the loco encountered so far was an AWS failure on Tuesday. Angie. Mike B and I were at Grosmont when the loco returned to the shed so we were able to carry out a repair and clean the wheels, frames and rods while we were on. Our 22.00 finish was a little later than planned but it was worth the effort. Other than that all has been satisfactory.

The loco is due to work the Grosmont – Battersby service again tomorrow and, subject to confirmation, Grosmont – Whitby on Monday and Tuesday next week.

 

26th March 2021

Print

Following the last update we steamed the K1 on Wednesday with Angie, Rich P and me (for refitting the tender steam brake piston and our steam/functional test) and today with Angie and me (for fitting new piston packings (kindly collected from Steve Andrews at Derby by Mike B) and the annual mainline hot exam/functional test).

The tests on both days were satisfactory and, before heading home this evening, we persuaded the Doc (who had been messing about with Bill on the Q6) to drive the engine up and down the head shunt a few times. He must have been impressed as he gave Angie and me access to the Q6 team’s stash of custard creams afterwards – only one each mind.

The next working party on the K1 will be on Sunday when we will be preparing the loco for working the Grosmont – Battersby crew competence trains next week.

 

23rd March 2021

Print

Since the last report on the K1, working parties have taken place most days and several nights. Angie, Mike B, Chris H, Scott M, Kenny M, Rolly and me have all been involved

With the fire lit, the K1 stands on the pit under the coaling plant - Paul Hutchinson

In 35 years of working on the K1 I have never known it object so much to being put back together. It is however very nearly there. After a false start on Sunday the fire was lit late this afternoon for a NELPG steam/functional test tomorrow - we still need to refit the tender steam brake piston (after a replacement little end pin was fitted today) but that will be done first thing in the morning. The attached photo was taken as Angie and I left Grosmont at 17.30ish today - an early finish for once.

The mainline hot exam/functional test is scheduled to take place on Friday and the loco is due work crew competency trains to Battersby Monday to Thursday next week. If you would like to be involved in looking after the loco while it does this work then please let me know.

With regard to the coach, JP has made good progress with the new 24v lighting system, tidying up much of the 240v wiring at the same time. Richard De and Chris H have also sorted out a number of very sticky drop lights and Kenny M has started the alterations to the battery boxes necessary to accommodate the new batteries.

 

11th March 2021

Print

Since the last report there has been 5 working parties on the K1 (Scott M, Mike B, Angie, Charlie Woods, Chris Cubit and me all did at least one day). In addition JP has done another day working on the coach lighting.

The work undertaken was as follows:

  1. Both pistons (having been collected by Mike last Friday) refitted following final cleaning, dressing and fitting of rings.
  2. Piston packing followers, lense rings, springs and spring cups cleaned (both sides).
  3. RH piston rod reunited with crosshead.
  4. LH cylinder cover cleaned and prepared for refitting.
  5. LH crosshead shimmed.
  6. Driving wheelset hornstay bolt hole reaming completed, replacement bolts fitted and secured.
  7. Driving wheelset springs lifted into hangers.
  8. New cab lights ,including a lamp over the reverser indicator, fitted (by Charlie and Chris C).
  9. Machining new snifting valve completed (on my lathe at home).
  10. More filling and rubbing down carried out to repaired areas on the back of the tender.
  11. Rear draw gear removed. Draw bar (hook) die pen tested (ok) and delivered to Flavells for repair – from there it will go to Graythorpe Forge for normalising.
 

1st March 2021

Print

We had working parties on Wednesday and Thursday (Angie and me working on the K1 both days) and Saturday (Angie, Mike B, Scott M and me on the K1).

Over the 3 days the following work was undertaken:

  1. Annual service of blower valve completed and refitted.
  2. 5 hornstay bolt holes on driving wheel set reamed true and new bolts fitted to 4 of them. 3 more holes to ream and 4 bolts to fit.
  3. Reaming of tender brake piston and con rod little end holes started.
  4. Both cylinders cleaned internally, cover studs cleaned and sealing faces prepared for refitting.
  5. Both driver's side injector intermediate air ministry joints split (joints to be remade).
  6. Steam chest pressure gauge supply pipe annealed, front section refitted, union fittings removed from intermediate section - to be replaced with new fittings.
  7. Two cracks to the rear of tender tank welded and filled - further filling and painting to be progressed this week (weather permitting).
  8. Building up (with weld) of thin areas of the smokebox door sealing ring completed.
  9. New baffle plate trimmed to correct profile/shape.
 

24th February 2021

Print

Steady progress has been made with the K1 with working parties on Thursday (ex junior Scott Middlemiss, Angie and me) and Satuday (Scott, Rolly (who drove over from Morecambe) Mike B and me). The following work was undertaken over the two days:

  1. Both driving wheel set springs removed.
  2. Driving wheelset spring hangers, hanger bolts, Spencer blocks and hornstay areas thoroughly cleaned.
  3. All driving wheelset hornstay bolt split pins removed.
  4. Weld repairs to smokebox door seal ring started (to complete this week).
  5. Annual service of blower valve started (to complete this week).
  6. R.T. under keep tray drain pipework reassembled and braised in to position following weld repair to the tray (by Flavells). Tray refitted.
  7. Both little end pins die pen tested (both ok).
  8. Drivers side injector overflow pipe removed, fractured braise to elbow/flange repaired and pipe refitted.
  9. Steam chest pressure gauge feed pipe removed for repair (frost damage). Repair and refitting to be carried out this week.
  10. Primary screen removed from smoke box.
  11. Fire hole bottom protector plate removed.
  12. Drivers side injector pipework clips and brackets removed ahead of remaking the intermediate air ministry joints.

In addition to the above, JP joined us on Thursday afternoon to progress the electrical work on the coach.

 

Tuesday 18th February 2021

Print

We had three days working on the K1 last week and one day on the J27

Thursday’s working party was Mike B, Angie and me, Friday’s was Chris H, Angie and me and we were joined in the afternoon by JP who came to progress the new lighting system in the coach. Other then JP’s work on the coach both days were spent on the K1.  Friday and Saturday in particular were very cold – if it got above freezing it did not feel like it!

The work on the locos undertaken over the three days was:

  1. Both crossheads/pistons split
  2. Pistons removed, cleaned and rings removed.
  3. Pistons loaded into company van – piston rods to be ground true.
  4. Both trailing loco brake pull rods removed, cleaned and loaded into my van (they won’t fit in the company van). The adjuster threads are to be refurbished and one bush is to be replaced.
  5. Annual service completed to drain cocks.
  6. Smokebox door seal removed and channel cleaned.
  7. Tender brake piston removed - Con rod little end pin holes to be bushed.
  8. Crosshead clearances checked
  9. Leading side rods moved from Dev shed to NYMR machine shop where the leading bushes are to be replaced by the NYMR.

Some of these jobs took longer than usual because of the extreme cold. The oxy gear was in regular use gently warming many components up before they could be safely worked upon. Even so, when the pistons were removed the rings were frozen into the groves and the piston packings were full of ice as were the drain cock valves.

 

Saturday 6th February 2021

Print

A small working party of six carried out a few jobs on the J27 and K1 on Saturday. At the same time an even smaller working party of one made a start on sorting the coach lighting out.

62005 back in Deviation Shed on 4 February 2021 - Ian Pearson

Dr Pearson (the split pin king) took our temperatures to start with, following which the cylinder covers were removed from the J27 and the side rods and cylinder covers were removed from the K1 with all side rod bush clearances being checked as we went along. The K1's tender tank was also drained.

In addition the long awaited lighting upgrade was started on the coach. This is being carried out by John Paul Venus (JP) who volunteered to do the job in a moment of weakness. In short both the 240v and 24v lighting systems, including all existing wiring, are to be largely removed and replaced with a new 24v system with modern LED fittings. The batteries are also being replaced as is the charger. The new one will be a conditioning charger which will maintain the batteries rather than boiling them dry.

The 240v sockets will be left as they are for now but we are considering fitting an inverter for charging mobile phones etc while we are on the move.

When JP has finished the electrical work we aim to replace the roof lining and carry out a few joinery repairs in the coach kitchen before repainting the kitchen walls and floor. While I don't expect to be trampled in the rush, anyone wanting to get involved in this work would be more than welcome.

 
  • «
  •  Start 
  •  Prev 
  •  1 
  •  2 
  •  3 
  •  4 
  •  5 
  •  6 
  •  7 
  •  8 
  •  9 
  •  10 
  •  Next 
  •  End 
  • »


Page 1 of 15